# ManiGaussian: Dynamic Gaussian Splatting for Multi-task Robotic Manipulation

### v1

@@ Abstract
Performing language-conditioned robotic manipulation tasks in unstructured environments is highly demanded for general intelligent robots. Conventional robotic manipulation methods usually learn semantic representation of the observation for action prediction, which ignores the scene-level spatiotemporal dynamics for human goal completion. In this paper, we propose a dynamic Gaussian Splatting method named ManiGaussian for multi-task robotic manipulation, which mines scene dynamics via future scene reconstruction. Specifically, we first formulate the dynamic Gaussian Splatting framework that infers the semantics propagation in the Gaussian embedding space, where the semantic representation is leveraged to predict the optimal robot action. Then, we build a Gaussian world model to parameterize the distribution in our dynamic Gaussian Splatting framework, which provides informative supervision in the interactive environment via future scene reconstruction. We evaluate our ManiGaussian on RLBench tasks with variations, and the results demonstrate our framework can outperform the state-of-the-art methods by % in average success rate 111Project page: https://guanxinglu.github.io/ManiGaussian/.
@@
Designing autonomous agents for language-conditioned manipulation tasks [28, 2, 26, 70, 54, 55, 9] has been highly desired in the pursuit of artificial intelligence for a long time. In realistic deployment, intelligent robots are usually required to deal with unseen scenarios in novel tasks. Therefore, comprehending complex 3D structures in the deployment scenes is necessary for the robots to achieve high task success rates across diverse manipulation tasks.
To address the challenges, previous arts have made great processes in general manipulation policy learning, which can be divided into two categories including perceptive methods and generative methods. For the first regard, semantic features extracted by perceptive models are directly leveraged to predict the robot actions according to the visual input such as image [36, 13, 12], point cloud [5, 11, 71] and voxel [25, 55]. However, the perceptive methods heavily rely on multi-view cameras to cover the whole workbench to deal with the occlusion problem within unstructured environments, which restricts their deployment. To this end, generative methods [43, 42, 47, 31, 20, 68, 69, 32] capture the 3D scene structure information by reconstructing the scene and objects in arbitrary novel views with self-supervised learning. Nevertheless, they ignore the spatiotemporal dynamics that depict the physical interaction among objects during manipulation, and the predicted actions still fail to complete human goals without correct object interactions. Figure 1 shows a comparison of manipulation achieved by the conventional generative manipulation method (top) and the proposed method (bottom), where the conventional method fails to stack the two rose blocks due to the poor comprehension of scene dynamics.
@@ 2 Related Work
Visual Representations for Robotic Manipulation. Developing intelligent agents for language-conditioned manipulation tasks in complex and unstructured environments has been a longstanding objective. One of the key bottlenecks in achieving this goal is effectively representing visual information of the scene. Prior arts can be categorized into two branches: perceptive methods and generative methods. Perceptive methods directly utilize pretrained 2D [36, 13, 12, 71] or 3D visual representation backbone [5, 11, 25, 55] to learn scene embedding, where optimal robot actions are predicted based on the scene semantics. For example, InstructRL [36] and Hiveformer [13] directly passed 2D visual tokens through a multi-modal transformer to decode gripper actions, but struggled to handle complex manipulation tasks due to the lack of geometric understanding. To incorporate 3D information beyond images, PolarNet [5] and Act3D [11] utilized point cloud representation, where PolarNet used a PointNeXt [44]-based architecture and Act3D designed a ghost point sampling mechanism to decode actions. Moreover, PerAct [55] fed voxel tokens into a PerceiverIO [23]-based transformer policy, demonstrating impressive performance in a variety of manipulation tasks. However, perceptive methods heavily relied on seamless camera overlay for comprehensive 3D understanding, which makes them less effective in unstructured environments. To address this, generative methods [43, 42, 47, 31, 20, 68, 69, 32] have gained attention. which learns the 3D geometry through self-supervised novel view reconstruction. For instance, Li et al. [32] combined NeRF and time contrastive learning to embed 3D geometry and learn fluid dynamics within an autoencoder framework. GNFactor [69] optimized a generalizable NeRF with a reconstruction loss besides behavior cloning, and showed effective improvement in both simulated and real scenarios. However, conventional generative methods usually ignore the scene-level spatiotemporal dynamics that demonstrate the interaction among objects, and the predicted actions still fail to achieve human goals because of incorrect interaction.
@@
World Models. In recent years, world models have emerged as an effective approach to encode scene dynamics by predicting the future states given the current state and actions, which are explored in autonomous driving [57, 10, 21, 22], game agent [14, 15, 16, 17, 18, 66, 49] and robotic manipulation [19, 61, 50]. Early works [14, 15, 16, 17, 18, 19, 66, 49, 50] learned a latent space for future prediction by autoencoding, which acquired notable effectiveness in both simulated and real-world situations [61]. However, learning latent for accurate future prediction requires large amount of data and is limited to simple tasks such as robot control due to the weak representative ability of the implicit features. To address these limitations, explicit representation in the image domain [7, 51, 60, 40] and the language domain [34, 38, 58] has been widely studied because of the rich semantics. UniPi [7] reconstructed the future images with a text-conditional video generation model, employing an inverse dynamics model to obtain the intermediate actions. Dynalang [34] learned to predict text representations as future states, and enabled embodied agents to navigate in photorealistic home scans under human instructions. In contrast to these approaches, we generalize the world model to embedding space of dynamic Gaussian Splatting, which predicts the future state for the agent to learn scene-level dynamics from interactive environments.
@@
Gaussian Splatting. Gaussian Splatting [29] models the scenes with a set of 3D Gaussians which are projected to 2D planes with efficient differentiable splatting. Gaussian Splatting achieves higher effectiveness and efficiency compared with implicit representations such as Neural Radiance Fields (NeRF) [41, 27, 35, 32, 6, 53, 69] with fast inference, high fidelity, and strong editability for novel view synthesis. To deploy Gaussian Splatting in diverse complex scenarios, many variants have been proposed to enhance the generalization ability, enrich the semantic information, and reconstruct deformable scenes. For higher generalization ability across diverse scenes, recent works [56, 72, 74, 4, 8, 63] constructed a direct mapping from pixels to Gaussian parameters, where the latent features were learned from large-scale datasets. To integrate rich semantic information into Gaussian Splatting, many efforts [45, 75, 52] have been demonstrated in distilling Gaussian radiance fields from pretrained foundation models [46, 3, 48]. For instance, LangSplat [45] advanced the Gaussian representation by encoding language features distilled from CLIP [46] using a scene-wise language autoencoder, enabling efficient open-vocabulary localization compared with its NeRF-based counterpart [30]. For deformation modeling, time-variant Gaussian radiance fields [64, 62, 65, 39, 59, 33, 1] were reconstructed from videos instead of images, which are widely applied in applications such as surgical scene reconstruction [73, 37]. Although these approaches have achieved high-quality reconstruction from entire videos like interpolation, extrapolation to future states conditioned on previous states and actions is unexplored, which holds significance for scene-level dynamics modeling for interactive agents. In this paper, we formulate a dynamic Gaussian Splatting framework to model the scene dynamics of objects interactions, which enhance the physical reasoning for agents to complete a wide range of robotic manipulation tasks.
@@ 3 Approach
In this section, we first briefly introduce preliminaries on the problem formulation (Section 3.1), and then we present an overview of our pipeline (Section 3.2). Subsequently, we introduce our dynamic Gaussian Splatting framework (Section 3.3) that infers the semantics propagation of the manipulation scenarios in the Gaussian embedding space. To enable our dynamic Gaussian Splatting framework to learn scene dynamics from the interactive environment, we build a Gaussian world model (Section 3.4) that reconstructs future scenes according to the propagated semantics.
@@ 3.1 Problem Formulation
The demand for language-conditioned robotic manipulation is a significant aspect in the development of general intelligent robots. The agent is required to interactively predict the subsequent pose of the robot arm based on the observation and achieve the pose with a low-level motion planner to complete a wide range of manipulation tasks described in humans. The visual input at the step for the agent is defined as , where and respectively represent the single-view images and the depth images. The proprioception matrix indicates the gripper state including the end-effector position, openness, and current timestep. Based on the visual input and the language instructions, the agent is required to generate the optimal action for the robot arm and grippers , which respectively demonstrates the modification of the robot arm in translation, rotation, the motion-planner state and of the grippers in open state. To learn the manipulation policy effectively, expert demonstrations as offline datasets are provided for imitation learning, where the sample triplets contain the visual input, language instruction and the expert actions. Existing methods leverage powerful visual representation to learn informative latent features for optimal action prediction. However, they ignore the spatiotemporal dynamics which depicts the physical interaction among objects, and the predicted actions usually fail to complete complex human goals without correct object interactions. On the contrary, we present a dynamic Gaussian Splatting framework to mine the scene dynamics for robotic manipulation.
@@ 3.2 Overall Pipeline
The overall pipeline of our ManiGaussian method is shown in Figure 2, in which we construct a dynamic Gaussian Splatting framework that models the propagation of diverse semantic features in the Gaussian embedding space for manipulation. We also build a Gaussian world model to parameterize distributions in our dynamic Gaussian Splatting framework, which can provide informative supervision of scene dynamics by future scene reconstruction. More specifically, we transform the visual input from RGB-D cameras to a volumetric representation by lifting and voxelization for data preprocessing. For dynamic Gaussian Splatting, we leverage a Gaussian regressor to infer the Gaussian distribution of geometric and semantic features in the scene, which are propagated along time steps with rich scene-level spatiotemporal dynamics. For the Gaussian world model, we instantiate a deformation field to reconstruct the future scene according to the current scene and the robot actions, and require consistency between reconstructed and realistic scenes for dynamics mining. Therefore, the spatiotemporal dynamics indicating object correlation can be embedded into the geometric and semantic features learned in the dynamic Gaussian Splatting framework. Finally, we employ multi-modal transformer PerceiverIO [23] to predict the optimal robot actions for general manipulation tasks, which considers geometric and semantic features in Gaussian embedding space with human language instructions.
@@ 3.3 Dynamic Gaussian Splatting for Robotic Manipulation
In order to capture the scene-level dynamics for general manipulation tasks, we propose a dynamic Gaussian Splatting framework that models the propagation of diverse semantic features within the Gaussian embedding space. While the vanilla Gaussian Splatting has remarkable effectiveness and efficiency in reconstructing static environments, it fails to capture the scene dynamics for manipulation due to the lack of temporal information. To this end, we formulate a dynamic Gaussian Splatting framework based on the vanilla Gaussian Splatting methodology by enabling the Gaussian points of scene representation to move with robotic manipulation, which demonstrates the physical interactions between objects. The scene representation of our dynamic Gaussian Splatting framework contains the geometric features depicting the explicit visual clues and the semantic features illustrating the implicit high-level visual features, which are utilized to predict the optimal action for the robot arm and grippers.
@@
Dynamic Gaussian Splatting. Gaussian Splatting [29] is a promising approach for multi-view 3D reconstruction, which exhibits fast inference, high fidelity, and strong editability of generated content compared with Neural Radiance Field (NeRF) [41]. Gaussian Splatting represents a 3D scene explicitly with multiple Gaussian primitives, where the Gaussian primitive is parameterized by , where respectively represent the positions, rotation, color, scale, and opacity for the Gaussian primitive. To render a novel view, we project Gaussian primitives onto the 2D plane by differential tile-based rasterization. The value of the pixel can be rendered by the alpha-blend rendering:
@@
where is the rendered image, denotes the number of Gaussians in this tile, represents the 2D density of the Gaussian points in the splatting process, and stands for the covariance matrix acquired from the rotation and scales of the Gaussian parameters. However, the vanilla Gaussian Splatting encounters difficulties in reconstructing temporal information in changing environments, which limits the ability to model scene-level dynamics that is crucial for manipulation tasks. To address this, we enable the Gaussian particles to be propagated with time to capture the spatiotemporal dynamics of the scene. The parameters of the Gaussian primitive at the step can be expressed as follows:
@@
The positions, rotation, colors, scale and opacity with the superscript represent their counterparts at the step in the propagation, and is the high-level semantic features distilled from the Stable Diffusion [48] visual encoder based on the RGB images of the scene. In robotic manipulation, all objects are regarded as rigid body without inherent properties including colors, scales, opacity and semantic features, and , , and are therefore regarded as a time-independent parameters. The positions and rotation for Gaussian particles change during the manipulation due to the physical interaction between objects and robot grippers, which can be formulated as follows:
@@
Gaussian World Model. In our implementation, we present a Gaussian world model to parameterize the Gaussian mixture distribution in dynamic Gaussian Splatting, through which the future scene can be reconstructed via the parameter propagation. Therefore, the dynamic Gaussian Splatting model can acquire informative supervision in the interactive environment by considering the consistency between the reconstructed and realistic feature scenes. World models are effective to learn the environmental dynamics for downstream tasks by anticipating the future state based on the current state and action at the step, which have been applied to a variety of tasks including autonomous driving [57, 10, 21, 22] and game agent [14, 15, 16, 17, 18, 66, 49]. For our robotic manipulation tasks, we instantiate the current state in the world model as the visual observation in current step, and actions refer to the those of the robot arm and grippers. They are leveraged to predict the visual scenes observed in the next step that represents the future state. More specifically, the Gaussian world model contains a representation network that learns high-level visual features with rich semantics for the input observation, a Gaussian regressor that predicts the Gaussian parameters of different primitives based on the visual features, a deformation predictor that infers the difference of Gaussian parameters during the propagation, and a Gaussian renderer that generates the RGB images for the predicted future state:

### v2

@@ Abstract
Performing language-conditioned robotic manipulation tasks in unstructured environments is highly demanded for general intelligent robots. Conventional robotic manipulation methods usually learn a semantic representation of the observation for action prediction, which ignores the scene-level spatiotemporal dynamics for human goal completion. In this paper, we propose a dynamic Gaussian Splatting method named ManiGaussian for multi-task robotic manipulation, which mines scene dynamics via future scene reconstruction. Specifically, we first formulate the dynamic Gaussian Splatting framework that infers the semantics propagation in the Gaussian embedding space, where the semantic representation is leveraged to predict the optimal robot action. Then, we build a Gaussian world model to parameterize the distribution in our dynamic Gaussian Splatting framework, which provides informative supervision in the interactive environment via future scene reconstruction. We evaluate our ManiGaussian on RLBench tasks with variations, and the results demonstrate our framework can outperform the state-of-the-art methods by % in average success rate111Project page: https://guanxinglu.github.io/ManiGaussian/.
@@
Designing autonomous agents for language-conditioned manipulation tasks [63, 62, 31, 2, 29, 77, 59, 60, 11] has been highly desired in the pursuit of artificial intelligence for a long time. In realistic deployment, intelligent robots are usually required to deal with unseen scenarios in novel tasks. Therefore, comprehending complex 3D structures in the deployment scenes is necessary for the robots to achieve high task success rates across diverse manipulation tasks.
To address the challenges, previous arts have made great progress in general manipulation policy learning, which can be divided into two categories including perceptive methods and generative methods. For the first regard, semantic features extracted by perceptive models are directly leveraged to predict the robot actions according to the visual input such as image [40, 15, 14], point cloud [7, 13, 79] and voxel [28, 60]. However, the perceptive methods heavily rely on multi-view or gripper-mounted cameras to cover the whole workbench to deal with the occlusion problem within unstructured environments, which restricts their deployment. To this end, generative methods [48, 47, 52, 35, 22, 75, 76, 36] capture the 3D scene structure information by reconstructing the scene and objects in arbitrary novel views with self-supervised learning. Nevertheless, they ignore the spatiotemporal dynamics that depict the physical interaction among objects during manipulation, and the predicted actions still fail to complete human goals without correct object interactions. Figure 1 shows a comparison of manipulation achieved by the conventional generative manipulation method (top) and the proposed method (bottom), where the conventional method fails to stack the two rose blocks due to the poor comprehension of scene dynamics.
@@ 2 Related Work
Visual Representations for Robotic Manipulation. Developing intelligent agents for language-conditioned manipulation tasks in complex and unstructured environments has been a longstanding objective. One of the key bottlenecks in achieving this goal is effectively representing visual information of the scene. Prior arts can be categorized into two branches: perceptive methods and generative methods. Perceptive methods directly utilize pretrained 2D [40, 15, 14, 79] or 3D visual representation backbone [7, 13, 28, 60] to learn scene embedding, where optimal robot actions are predicted based on the scene semantics. For example, InstructRL [40] and Hiveformer [15] directly passed 2D visual tokens through a multi-modal transformer to decode gripper actions, but struggled to handle complex manipulation tasks due to the lack of geometric understanding. To incorporate 3D information beyond images, PolarNet [7] and Act3D [13] utilized point cloud representation, where PolarNet used a PointNeXt [49]-based architecture and Act3D designed a ghost point sampling mechanism to decode actions. Moreover, PerAct [60] fed voxel tokens into a PerceiverIO [26]-based transformer policy, demonstrating impressive performance in a variety of manipulation tasks. However, perceptive methods heavily rely on seamless camera overlay for comprehensive 3D understanding, which makes them less effective in unstructured environments. To address this, generative methods [48, 47, 52, 35, 22, 75, 76, 36] have gained attention. which learns the 3D geometry through self-supervised novel view reconstruction. For instance, Li et al. [36] combined NeRF and time contrastive learning to embed 3D geometry and learn fluid dynamics within an autoencoder framework. GNFactor [76] optimized a generalizable NeRF with a reconstruction loss besides behavior cloning, and showed effective improvement in both simulated and real scenarios. However, conventional generative methods usually ignore the scene-level spatiotemporal dynamics that demonstrate the interaction among objects, and the predicted actions still fail to achieve human goals because of incorrect interaction.
@@
World Models. In recent years, world models have emerged as an effective approach to encode scene dynamics by predicting the future states given the current state and actions, which are explored in autonomous driving [64, 12, 24, 25], game agent [16, 17, 18, 19, 20, 73, 54] and robotic manipulation [21, 68, 55]. Early works [16, 17, 18, 19, 20, 21, 73, 54, 55] learned a latent space for future prediction by autoencoding, which acquired notable effectiveness in both simulated and real-world situations [68]. However, learning latent for accurate future prediction requires a large amount of data and is limited to simple tasks such as robot control due to the weak representative ability of the implicit features. To address these limitations, explicit representation in the image domain [9, 56, 67, 45] and the language domain [38, 43, 6, 65, 23] has been widely studied because of the rich semantics. UniPi [9] reconstructed the future images with a text-conditional video generation model, employing an inverse dynamics model to obtain the intermediate actions. Dynalang [38] learned to predict text representations as future states, and enabled embodied agents to navigate in photorealistic home scans under human instructions. In contrast to these approaches, we generalize the world model to embedding space of dynamic Gaussian Splatting, which predicts the future state for the agent to learn scene-level dynamics from interactive environments.
@@
Gaussian Splatting. Gaussian Splatting [32] models the scenes with a set of 3D Gaussians which are projected to 2D planes with efficient differentiable splatting. Gaussian Splatting achieves higher effectiveness and efficiency compared with implicit representations such as Neural Radiance Fields (NeRF) [46, 30, 39, 36, 8, 58, 76] with fast inference, high fidelity, and strong editability for novel view synthesis. Please refer to [5] for a comprehensive survey on 3D Gaussian Splatting. To deploy Gaussian Splatting in diverse complex scenarios, many variants have been proposed to enhance the generalization ability, enrich the semantic information, and reconstruct deformable scenes. For higher generalization ability across diverse scenes, recent works [61, 80, 83, 4, 10, 70, 78] constructed a direct mapping from pixels to Gaussian parameters from large-scale datasets. To integrate rich semantic information into Gaussian Splatting, many efforts [50, 84, 57, 81] have been demonstrated in distilling Gaussian radiance fields from pretrained foundation models [51, 3, 53, 34]. For instance, LangSplat [50] advanced the Gaussian representation by encoding language features distilled from CLIP [51] using a scene-wise language autoencoder, enabling efficient open-vocabulary localization compared with its NeRF-based counterpart [33]. For deformation modeling, time-variant Gaussian radiance fields [71, 69, 72, 44, 66, 37, 1] were reconstructed from videos instead of images, which are widely applied in applications such as surgical scene reconstruction [82, 42]. Although these approaches have achieved high-quality reconstruction from entire videos like interpolation, extrapolation to future states conditioned on previous states and actions is unexplored, which holds significance for scene-level dynamics modeling for interactive agents. In this paper, we formulate a dynamic Gaussian Splatting framework to model the scene dynamics of object interactions, which enhances the physical reasoning for agents to complete a wide range of robotic manipulation tasks.
@@ 3 Approach
In this section, we first briefly introduce preliminaries on the problem formulation (Section 3.1), and then we present an overview of our pipeline (Section 3.2). Subsequently, we introduce a dynamic Gaussian Splatting framework (Section 3.3) that infers the propagation semantics of the manipulation scenarios in the Gaussian embedding space. To enable our dynamic Gaussian Splatting framework to learn scene dynamics from the interactive environment, we build a Gaussian world model (Section 3.4) that reconstructs future scenes according to the propagated semantics.
@@ 3.1 Problem Formulation
The demand for language-conditioned robotic manipulation is a significant aspect in the development of general intelligent robots. The agent is required to interactively predict the subsequent pose of the robot arm based on the observation and achieve the pose with a low-level motion planner to complete a wide range of manipulation tasks described in humans. The visual input at the step for the agent is defined as , where and respectively represent the single-view images and the depth images. The proprioception matrix indicates the gripper state including the end-effector position, openness, and current timestep. Based on the visual input and the language instructions, the agent is required to generate the optimal action for the robot arm and grippers , which respectively demonstrates the target translation in voxel , rotation , openness and collision avoidance .
To learn the manipulation policy effectively, expert demonstrations as offline datasets are provided for imitation learning, where the sample triplets contain the visual input, language instruction and expert actions. Existing methods leverage powerful visual representations to learn informative latent features for optimal action prediction. However, they ignore the spatiotemporal dynamics which depicts the physical interaction among objects, and the predicted actions usually fail to complete complex human goals without correct object interactions. On the contrary, we present a dynamic Gaussian Splatting framework to mine the scene dynamics for robotic manipulation.
@@ 3.2 Overall Pipeline
The overall pipeline of our ManiGaussian method is shown in Figure 2, in which we construct a dynamic Gaussian Splatting framework that models the propagation of diverse semantic features in the Gaussian embedding space for manipulation. We also build a Gaussian world model to parameterize distributions in our dynamic Gaussian Splatting framework, which can provide informative supervision of scene dynamics by future scene reconstruction. More specifically, we transform the visual input from RGB-D cameras to a volumetric representation by lifting and voxelization for data preprocessing. For dynamic Gaussian Splatting, we leverage a Gaussian regressor to infer the Gaussian distribution of geometric and semantic features in the scene based on the representation, propagated along time steps with rich scene-level spatiotemporal dynamics. For the Gaussian world model, we instantiate a deformation field to reconstruct the future scene according to the current scene and the robot actions, and require consistency between reconstructed and realistic scenes for dynamics mining. Therefore, the spatiotemporal dynamics indicating object correlation can be embedded into the representation learned in the dynamic Gaussian Splatting framework. Finally, we employ multi-modal transformer PerceiverIO [26] to predict the optimal robot actions for general manipulation tasks, which considers geometric, semantic, and dynamic information with human instructions.
@@ 3.3 Dynamic Gaussian Splatting for Robotic Manipulation
In order to capture the scene-level dynamics for general manipulation tasks, we propose a dynamic Gaussian Splatting framework that models the propagation of diverse semantic features within the Gaussian embedding space. While the vanilla Gaussian Splatting has remarkable effectiveness and efficiency in reconstructing static environments, it fails to capture the scene dynamics for manipulation due to the lack of temporal information. To this end, we formulate a dynamic Gaussian Splatting framework based on the vanilla Gaussian Splatting methodology by enabling the Gaussian points of scene representation to move with robotic manipulation, which demonstrates the physical interactions between objects. The scene representation contains geometric information depicting the explicit visual clues, semantic information illustrating the implicit high-level visual features, and dynamic information encoding the physical properties of the scene, which are utilized to predict the optimal actions.
@@
Dynamic Gaussian Splatting. Gaussian Splatting [32] is a promising approach for multi-view 3D reconstruction, which exhibits fast inference, high fidelity, and strong editability of generated content compared with Neural Radiance Field (NeRF) [46]. Gaussian Splatting represents a 3D scene explicitly with multiple Gaussian primitives, where the Gaussian primitive is parameterized by , where respectively represent the positions, color, rotation, scale, and opacity for the Gaussian primitive. To render a novel view, we project Gaussian primitives onto the 2D plane by differential tile-based rasterization. The value of the pixel can be rendered by the alpha-blend rendering:
@@
where is the rendered image, denotes the number of Gaussians in this tile, represents the 2D density of the Gaussian points in the splatting process, and stands for the covariance matrix acquired from the rotation and scales of the Gaussian parameters. However, the vanilla Gaussian Splatting encounters difficulties in reconstructing changing environments, which limits the ability to model scene-level dynamics that is crucial for manipulation tasks. To address this, we enable the Gaussian particles to be propagated with time to capture the spatiotemporal dynamics of the scene. The parameters of the Gaussian primitive at the step can be expressed as follows:
@@
The positions, colors, rotations, scales, and opacities with the superscript represent their counterparts at the step in the propagation, and is the high-level semantic feature distilled from the Stable Diffusion [53] visual encoder based on the RGB images of the scene. In robotic manipulation, all objects are regarded as rigid bodies without inherent properties including colors, scales, opacities, and semantic features. , , and are therefore regarded as time-independent parameters. The positions and rotation of Gaussian particles change during the manipulation due to the physical interaction between objects and robot grippers, which can be formulated as follows:
@@
Gaussian World Model. In our implementation, we present a Gaussian world model to parameterize the Gaussian mixture distribution in dynamic Gaussian Splatting, through which the future scene can be reconstructed via parameter propagation. Therefore, the dynamic Gaussian Splatting model can acquire informative supervision in the interactive environment by considering the consistency between the reconstructed and realistic feature scenes. World models are effective to learn the environmental dynamics for downstream tasks by anticipating the future state based on the current state and action at the step, which have been applied to a variety of tasks including autonomous driving [64, 12, 24, 25] and game agent [16, 17, 18, 19, 20, 73, 54]. For our robotic manipulation tasks, we instantiate the current state in the world model as the visual observation in the current step, and actions refer to those of the robot arm and grippers. They are leveraged to predict the visual scenes observed in the next step that represent the future state. More specifically, the Gaussian world model contains a representation network that learns high-level visual features with rich semantics for the input observation, a Gaussian regressor that predicts the Gaussian parameters of different primitives based on the visual features, a deformation predictor that infers the difference of Gaussian parameters during the propagation, and a Gaussian renderer (1) that generates the RGB images for the predicted future state: